OpenStudy (anonymous):

Medal if someone helps me understand this!!!!! Look at the following expression. Are there any values of x for which the expression is not defined under the set of real numbers? Explain your reasoning. 4/sqrtx+2

OpenStudy (anonymous):

\[\frac{ 4 }{ \sqrt{x+2} }\]

OpenStudy (amistre64):

there are a few major rules in math that must not be broken ... one of them is dividing by zero

OpenStudy (amistre64):

another major rule is, you CANNOT have a real solution to the sqrt of a negative number

OpenStudy (anonymous):

So the only number i could use is 0?

OpenStudy (amistre64):

what we need for this to be valid is the underside part to conform to the rules: it cant equal 0, and the innards have to remian postivie

OpenStudy (anonymous):

yeah that i know but its asking me to plug in a number that is not real as x

OpenStudy (amistre64):

x+2 > 0 has to fit

OpenStudy (amistre64):

a number that aint real?

OpenStudy (amistre64):

Are there any values of x for which the expression is not defined under the set of real numbers? its asking for the domain or real

OpenStudy (anonymous):

Yeah so either 0 or an imaginary number such as 5i

OpenStudy (amistre64):

*of real

OpenStudy (amistre64):

its asking to restrict your results to real solution ...

OpenStudy (anonymous):

So there really arent any number to plug in that are nureal then?

OpenStudy (anonymous):

unreal

OpenStudy (amistre64):

in other words, sqrt(-4) is not a real value :)

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

ok i understand! thanks so much.

OpenStudy (amistre64):

your welcome

OpenStudy (anonymous):

If x+2 <= 0 the world implodes